# LoRePIE: $\ell_{0}$ Regularised Extended Ptychographical Iterative Engine

### v1

@@ I Introduction
The use of a defocused probe at the sample plane enables scanning of a larger Field of View (FoV) with a reasonable acquisition time by increasing the scan step size, subject to a minimum overlap constraint between probe positions [^1] The 4-D STEM dataset of Rotavirus [8] used in this paper, which was acquired in approximately $16$s, yields a FoV of approximately $400{\rm nm}\times 400{\rm nm}$.. Phase information can be retrieved by using iterative methods [27, 24, 28, 22], given sufficient overlap between adjacent illuminated areas [29]. The electron fluence is reduced on account of the increased scan step size and larger probe size. In this work, we centre our attention on defocused ptychography discussed by Zhou et al. [8], which has been demonstrated as a suitable method for imaging extremely radiation sensitive biological samples.
@@
In this paper [^2] Preliminary results of this work were presented in [30, 31]., we introduce an $\ell_{0}$ regularisation of the ePIE algorithm [22], or in short LoRePIE . The $\ell_{0}$-quasi-norm [^3] In strict mathematical terms, $\ell_{0}$ defined as $\|\boldsymbol{u}\|_{0}\coloneqq|\{i:|u_{i}|\neq 0\}|$, which only counts the number of non-zero elements, is not a norm nor a pseudo-norm. regularisation, enforces the estimation of the object wave function to have a sparse representation – i.e., only a few coefficients are non-zero – in a transform domain, e.g., in a Discrete Cosine Transform (DCT) [32], Discrete Fourier Transform (DFT), or a Discrete Wavelet Transform (DWT) [33]. The underlying mechanism in $\ell_{0}$ regularisation is that natural images, e.g., those of interest here, can be distinguished from noise-like images based on their sparsity level in a transform domain. This principle has been leveraged by a large number of denoising and signal recovery algorithms [34, 35] and in the theory of compressive sensing [36, 37]. In the context of this paper, $\ell_{0}$ regularisation of the estimated object wave function reduces to only three computationally scalable steps: (i) transforming the phase and amplitude of the estimated object wave function to a suitable domain, which is a matter of choice or design, (ii) an application of a hard-thresholding operator with parameter $K$, which maps all but $K$ (absolute value) largest coefficients to zero, and (iii) transforming the processed coefficients back to the original domain. By artificially down-sampling the probe positions on experimental 4-D STEM data of double layered Rotavirus particles acquired with 85% overlap, we show LoRePIE is more robust to a reduced overlap ratio compared to conventional ePIE. These suggest that using this approach, the fluence delivered to the sample could be reduced by scanning larger areas without compromising the quality of the reconstructed object wave function compared.
@@
Assuming that data is collected in the far-field [^4] We note that the LoRePIE algorithm proposed in this paper can be directly applied to ptychography in the near-field or Fresnel regime [38]. or Fraunhofer plane, the wave function at the detector plane, can be modelled as the Fourier transform of the exit wave [18, Ch. 17], i.e.,
@@ III Ptychography methods: from ePIE to regularised ePIE
Our regularised algorithm is based on the ePIE algorithm. Therefore, in this section, we first provide a suitable mathematical formulation of ePIE followed by an adaptation that yields its $\ell_{0}$ regularised version.
@@ 4             $\boldsymbol{o}^{(t)}_{l}\leftarrow\boldsymbol{\Pi}_{l}\boldsymbol{o}^{(t)}$ // Cropped object FoV
We now introduce $\ell_{0}$ regularised ePIE that is robust to low probe overlap ratios. The underlying idea of LoRePIE is to regularise the estimation of object crops $\boldsymbol{o}^{(t)}_{l}$ assuming that both its phase and amplitude have a sparse representation in a transform domain. See Sec. S7-B for an illustration of this idea. As outlined in Algorithm 1, LoRePIE involves the same steps as those in ePIE algorithm, with an additional step (Line 11) that regularises the object data using a function $\mathcal{D}:\mathbb{C}^{N_{\rm d}}\mapsto\mathbb{C}^{N_{\rm d}}$.
@@
In comparison to ePIE, LoRePIE requires four design parameters: two regularisation parameters and two sparsity bases. The regularisation parameters $K_{\rm amp}$ and $K_{\rm phs}$ control the weight of the regularisations. When $K_{\rm amp}=K_{\rm phs}=N_{\rm d}$, LoRePIE reduces to ePIE, i.e., no regularisation. Low quality measured data requires stronger regularisation or, equivalently, smaller values for $K_{\rm amp}$ and $K_{\rm phs}$. The choice of the sparsity matrix depends on the application and the type of the feature present in the sample. For example, DCT and DFT are reasonable choices for samples with periodic features [33].
@@ IV Numerical results
Evaluation. To evaluate the performance of LoRePIE against ePIE and to avoid bias during comparison of the reconstructions generated, the reference object and probe data were obtained from 100 iterations of the ePIE algorithm followed by 100 iterations of the Difference Map (DM) [24] algorithm utilising the Ptypy [46] package. Details and illustrations of the reference data can be found in Secs. S7-A and  S7-B, which provide insight into why $\ell_{0}$-denoising enables the LoRePIE algorithm to achieve a higher quality estimate of the wave function. The remaining reconstructions were carried out using codes developed in MATLAB and available from [8].
Ablation studies. The hyperparameters were fine tuned as described in the Supplementary Information S7-CS7-D, and  S7-E. Various values of the regularisation parameters are tested in Sec. S7-C, which shows that setting $K_{\rm amp}=K_{\rm phs}=0.05N_{\rm d}$ yields the highest NRMSE values for all down-sampling factors. Based on our observations in Sec. S7-D, using 50 iterations of the LoRePIE and ePIE algorithms offers a reasonable trade-off between reconstruction quality and computational time. For the range of probe and object update step size parameters given in Sec. S7-E, we investigated efficient values of these parameters for the ePIE and LoRePIE algorithms for each down-sampling factor. For all subsequent reconstructions, we use the update step size values that yields the highest Structural Similarity Index Measure (SSIM)  [47] values in Fig. S15.
@@ VI Acknowledgement
S7-A Reference data
A reference data is needed to evaluate the performance of the proposed LoRePIE algorithm. The reference object and probe images are generated through an application of the ePIE [22] algorithm run for 100 iterations followed by the Difference Map (DM) [24] algorithm run for 100 iterations. Subsequent application of ptychographical algorithms is a typical routine performed, e.g., in [48, 49, 50], for avoiding stagnation of a phase retrieval process stuck at a local minimum. The LoRePIE algorithm used in this paper for electron ptychography builds upon the ePIE code developed in Matlab by the authors of [8]. Therefore, to mitigate bias in our analyses towards this implementation of ePIE and LoRePIE, we used Ptypy [46] that is a well established package for ptychography reconstructions in X-ray and electron microscopy communities. For the ePIE algorithm, the object and probe update step size parameters are set to $\alpha_{\rm o}=\alpha_{\rm p}=0.01$ following the notation in Lines 9 and 10 of Algorithm. 1. For the DM algorithm, both the object and probe inertia parameters are set to 0.001 and the mix parameter between DM and alternating projection is set to 0.75.
Fig. S4 shows the amplitude and phase components of the reference object and the amplitude of probe data. The Power Spectrum Density (PSD) of the object phase reveals periodic frequency features. By examining the line profile of the object phase in the PSD map, we note that these periodic features are due to the grid artefacts [51], since inverting the frequency of the first periodic peak in Fig. S4-bottom gives the scan step size: $1/0.32~{}{\rm nm}^{-1}=3.125~{}{\rm nm}$.
Figure S4: Reference data used for evaluation. These data are generated by 100 iterations of the ePIE algorithm followed by another 100 iterations of the DM algorithm. The periodic features in the PSD of the object data are due to the grid artefacts, since inverting the frequency of the first periodic peak in line profile curve gives the scan step size: $1/0.32~{}{\rm nm}^{-1}=3.125~{}{\rm nm}$.
S7-B Low-complexity of the Rotavirus phase image
@@
Figure S5: Low-complexity of the reference Rotavirus phase image. The relative error is computed between the reference phase image shown in Fig. S4 and its best $K_{\rm phs}$-term approximation in 2-D DCT domain. Keeping only 1% of the DCT coefficients and mapping the rest to zero results in an approximation of the reference phase image that is almost indistinguishable to human eyes with approximately 3.5% relative error.
S7-C Ablation study: impact of regularisation parameters on the performance of LoRePIE
@@
We consider down-sampling factor in the range of $\{1,2,3,4,5\}$, which is equivalent to subsampling ratios $\{1,1/4,1/9,1/16,1/25,1/36\}$. For each down-sampling factor, we obtain 121 reconstructions for every pair of the phase and amplitude regularisation parameters, i.e., $K_{\rm amp}/N_{\rm d}\in\{5,10,20,\cdots,100\}$ and $K_{\rm phs}/N_{\rm d}\in\{5,10,20,\cdots,100\}$. We compute the quality of every reconstructed phase image compared to the reference phase image shown in Fig. S4 in terms of the Structural Similarity Index Measure (SSIM) and the Peak Signal-to-Noise Ratio (PSNR). We also compute the Normalised Root Mean Squared Error (NRMSE) between the reference object and the complex-valued reconstructed object data.
@@
S7-D Ablation study: impact of number of iterations on the performance of LoRePIE and ePIE
Following the study in Sec. S7-C, we set $K_{\rm phs}=K_{\rm amp}=0.05N_{\rm d}$ and investigate the impact of the number of iterations on the performance of ePIE and LoRePIE algorithms.
@@
S7-E Ablation study: impact of update step size parameters on the performance of LoRePIE and ePIE
Based on our study in Secs. S7-C and S7-D, we set $K_{\rm phs}=K_{\rm amp}=0.05N_{\rm d}$ and set the number of ePIE and LoRePIE iterations to 50. In this section, we investigate the impact of the update step size parameters for object ($\alpha_{\rm o}$) and probe ($\alpha_{\rm p}$) in ePIE and LoRePIE algorithms.

### v2

@@ I Introduction
The use of a defocused probe at the sample plane enables scanning of a larger Field of View (FoV) with a reasonable acquisition time by increasing the scan step size, subject to a minimum overlap constraint between probe positions. For reference, the 4-D STEM dataset of Rotavirus [8] used in this paper, which was acquired in approximately $16$s, yields a FoV of approximately $400{\rm nm}\times 400{\rm nm}$. Phase information can be retrieved by using iterative methods [27, 24, 28, 22], given sufficient overlap between adjacent illuminated areas [29, 30]. The electron fluence is reduced on account of the increased scan step size and larger probe size. In this work, we centre our attention on defocused ptychography discussed by Zhou et al. [8], which has been demonstrated as a suitable method for imaging extremely radiation sensitive biological samples.
@@
In this paper, we introduce an $\ell_{0}$ regularisation of the ePIE algorithm [22], or in short LoRePIE . We note that in strict mathematical terms, $\ell_{0}$ defined as $\|\boldsymbol{u}\|_{0}\coloneqq|\{i:|u_{i}|\neq 0\}|$, which only counts the number of non-zero elements, is not a norm nor a pseudo-norm. The $\ell_{0}$-quasi-norm regularisation, enforces the estimation of the object wave function to have a sparse representation – i.e., only a few coefficients are non-zero – in a transform domain, e.g., in a Discrete Cosine Transform (DCT) [31], Discrete Fourier Transform (DFT), or a Discrete Wavelet Transform (DWT) [32]. The underlying mechanism in $\ell_{0}$ regularisation is that natural images, e.g., those of interest here, can be distinguished from noise-like images based on their sparsity level in a transform domain. This principle has been leveraged by a large number of denoising and signal recovery algorithms [33, 34] and in the theory of compressive sensing [35, 36]. In the context of this paper, $\ell_{0}$ regularisation of the estimated object wave function reduces to only three computationally scalable steps: (i) transforming the phase and amplitude of the estimated object wave function to a suitable domain, which is a matter of choice or design, (ii) an application of a hard-thresholding operator with parameter $K$, which maps all but $K$ (absolute value) largest coefficients to zero, and (iii) transforming the processed coefficients back to the original domain. By artificially down-sampling the probe positions on experimental 4-D STEM data of double layered Rotavirus particles acquired with 85% overlap, we show LoRePIE is more robust to a reduced overlap ratio compared to conventional ePIE. These suggest that using this approach, the fluence delivered to the sample could be reduced by scanning larger areas without compromising the quality of the reconstructed object wave function compared. Preliminary results of this work were presented in [37, 38].
@@
Assuming that data is collected in the far-field or Fraunhofer plane, the wave function at the detector plane, can be modelled as the Fourier transform of the exit wave [18, Ch. 17], i.e.,
@@ III Ptychography methods: from ePIE to regularised ePIE
Our regularised algorithm is based on the ePIE algorithm. Therefore, in this section, we first provide a suitable mathematical formulation of ePIE followed by an adaptation that yields its $\ell_{0}$ regularised version. We note that the proposed $\ell_{0}$ regularisation can be directly applied to ptychography in the near-field or Fresnel regime.
@@ 4             $\boldsymbol{o}^{(t)}_{l}\leftarrow\boldsymbol{\Pi}_{l}\boldsymbol{o}^{(t)}$ // Cropped object FoV
We now introduce $\ell_{0}$ regularised ePIE that is robust to low probe overlap ratios. The underlying idea of LoRePIE is to regularise the estimation of object crops $\boldsymbol{o}^{(t)}_{l}$ assuming that both its phase and amplitude have a sparse representation in a transform domain. See Sec. S1 in Supplementary Material for an illustration of this idea. As outlined in Algorithm 1, LoRePIE involves the same steps as those in ePIE algorithm, with an additional step (Line 11) that regularises the object data using a function $\mathcal{D}:\mathbb{C}^{N_{\rm d}}\mapsto\mathbb{C}^{N_{\rm d}}$.
@@
In comparison to ePIE, LoRePIE requires four design parameters: two regularisation parameters and two sparsity bases. The proposed regularisation modification to the ePIE algorithm does not cause divergence of the algorithm. However, the choice of regularisation parameters, $K_{\rm phs}$ and $K_{\rm amp}$, influences the quality of the recovered object. For an orthogonal sparsity basis, two extreme cases can be considered: (1) when $K_{\rm phs}=K_{\rm amp}=N_{\rm d}$, no regularisation is applied, and LoRePIE reduces to ePIE; (2) when $K_{\rm phs}=K_{\rm amp}=1$, LoRePIE estimates an object with constant amplitude and phase. For sufficiently large regularisation parameters, determined by the signal-to-noise ratio, LoRePIE outperforms ePIE.
The applicability of the proposed regularisation depends on the nature of the object and the choice of sparsity basis. Unlike total variation regularisation, our $\ell_{0}$ denoiser does not promote piecewise smoothness, making it less suitable for objects with highly structured features. Additionally, the choice of the sparsity domain significantly affects performance; for instance, enforcing sparsity in the DCT or DFT domains is beneficial for natural images with periodic features [32], but may not be optimal for highly oscillatory or amorphous structures. An additional consideration is parameter selection: excessively small values of $K_{\rm phs}$ and $K_{\rm amp}$ may suppress information, while large values reduce the regularisation effect. In principle, low quality, measured data requires stronger regularisation or, equivalently, smaller values for $K_{\rm amp}$ and $K_{\rm phs}$. Investigating alternative sparsity domains, such as DWT, and adaptive parameter selection strategies is ongoing.
@@
The choice of $\ell_{0}$ regularisation in our approach is motivated by its simplicity and effectiveness as a non-linear filtering technique. Unlike low-pass filters, which indiscriminately suppress high-frequency Fourier components, $\ell_{0}$ regularisation selects coefficients in a transform domain, Fourier or otherwise, based on their intensity, thereby preserving fine details while reducing noise. Applying $\ell_{0}$ regularisation separately to the amplitude and phase components of an object estimate offers flexibility in handling these independently. For instance, in the case of a phase-only object, a natural choice is to set $K_{\rm amp}=1$, which enforces a constant amplitude regardless of the regularisation of the phase component. This independent treatment of phase and amplitude also facilitates extensions to alternative regularisation approaches, such as enforcing sparsity in the gradient domain or replacing the $\ell_{0}$ quasi-norm with $\ell_{1}$ norm. These extensions are a focus of ongoing investigations.
Similar regularisations have been incorporated in ptychographic algorithms. For example, recently a half-quadratic splitting framework has been used to apply regularisations in the form of plug-and-play priors [43], and a quadratic regularisation for optical ptychography has been explored in [44] and reference [45] proposed a conjugate gradient descent algorithm regularised by an $\ell_{1}$ norm for electron ptychography. In the field of Fourier ptychography, where the object is illuminated using plane wave tilted illumination, as opposed to a convergent illumination in electron ptychography, regularisation techniques have been extensively applied in the object estimation process. Notable examples include total variation and Tikhonov [46], promoting $\ell_{0}$ sparsity in Block-Matching and 3D (BM3D) [47] frames [48], and image denoisers [49].
@@ IV Numerical results
Figure 2: Reference data used for evaluation. These data are generated by 100 iterations of the ePIE algorithm followed by another 100 iterations of the DM algorithm. The periodic features in the PSD of the object data are due to the grid artefacts, since inverting the frequency of the first periodic peak in line profile curve gives the scan step size: $1/0.32~{}{\rm nm}^{-1}=3.125~{}{\rm nm}$.
@@
Reference data. To evaluate the performance of LoRePIE against ePIE and to avoid bias during comparison of the reconstructions generated, the reference object and probe data were obtained from 100 iterations of the ePIE algorithm followed by 100 iterations of the Difference Map (DM) [24] algorithm utilising the well-established Ptypy [53] package. Subsequent application of ptychographical algorithms is a typical routine performed, e.g., in [54, 55, 56], for avoiding stagnation of a phase retrieval process stuck at a local minimum. For the ePIE algorithm, the object and probe update step size parameters were set to $\alpha_{\rm o}=\alpha_{\rm p}=0.01$. For the DM algorithm, both the object and probe inertia parameters are set to 0.001 and the mix parameter between DM and alternating projection is set to 0.75.
Fig. 2 shows the amplitude and phase components of the reference object and the amplitude of probe data. The Power Spectrum Density (PSD) of the object phase reveals periodic frequency features. By examining the line profile of the object phase in the PSD map, we note that these periodic features are due to the grid artefacts [57], since inverting the frequency of the first periodic peak in Fig. 2-bottom gives the scan step size: $1/0.32~{}{\rm nm}^{-1}=3.125~{}{\rm nm}$.
@@
Ablation studies. The hyperparameters were fine tuned as described in the Supplementary Material, Secs. S2, S3, and S4. Various values of the regularisation parameters are tested in Sec. S2, which shows that setting $K_{\rm amp}=K_{\rm phs}=0.05N_{\rm d}$ yields the highest NRMSE values for all down-sampling factors. Based on our observations in Sec. S3, using 50 iterations of the LoRePIE and ePIE algorithms offers a reasonable trade-off between reconstruction quality and computational time. For the range of probe and object update step size parameters given in Sec. S4, we investigated efficient values of these parameters for the ePIE and LoRePIE algorithms for each down-sampling factor. For all subsequent reconstructions, we use the update step size values that yields the highest Structural Similarity Index Measure (SSIM)  [58] values in Fig. S12.
@@ VI Acknowledgement
VI-A Funding
Rosalind Franklin Institute; UK Research and Innovation; Engineering and Physical Sciences Research Council; University of Liverpool; University of Oxford.
VI-B Disclosures
The authors declare no conflicts of interest.
VI-C Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
VI-D Supplemental document
See the supplementary material for ablation studies conducted to fine-tune the hyperparameters of the ePIE and LoRePIE algorithms.
@@ VI-D Supplemental document
S7-A Low-complexity of the Rotavirus phase image
@@
Figure S5: Low-complexity of the reference Rotavirus phase image. The relative error is computed between the reference phase image shown in Fig. 2 and its best $K_{\rm phs}$-term approximation in 2-D DCT domain. Keeping only 1% of the DCT coefficients and mapping the rest to zero results in an approximation of the reference phase image that is almost indistinguishable to human eyes with approximately 3.5% relative error.
S7-B Ablation study: impact of regularisation parameters on the performance of LoRePIE
@@
We consider down-sampling factor in the range of $\{1,2,3,4,5\}$, which is equivalent to subsampling ratios $\{1,1/4,1/9,1/16,1/25,1/36\}$. For each down-sampling factor, we obtain 121 reconstructions for every pair of the phase and amplitude regularisation parameters, i.e., $K_{\rm amp}/N_{\rm d}\in\{5,10,20,\cdots,100\}$ and $K_{\rm phs}/N_{\rm d}\in\{5,10,20,\cdots,100\}$. We compute the quality of every reconstructed phase image compared to the reference phase image shown in Fig. 2 in terms of the Structural Similarity Index Measure (SSIM) and the Peak Signal-to-Noise Ratio (PSNR). We also compute the Normalised Root Mean Squared Error (NRMSE) between the reference object and the complex-valued reconstructed object data.
@@
S7-C Ablation study: impact of number of iterations on the performance of LoRePIE and ePIE
Following the study in Sec. S7-B, we set $K_{\rm phs}=K_{\rm amp}=0.05N_{\rm d}$ and investigate the impact of the number of iterations on the performance of ePIE and LoRePIE algorithms.
@@
S7-D Ablation study: impact of update step size parameters on the performance of LoRePIE and ePIE
Based on our study in Secs. S7-B and S7-C, we set $K_{\rm phs}=K_{\rm amp}=0.05N_{\rm d}$ and set the number of ePIE and LoRePIE iterations to 50. In this section, we investigate the impact of the update step size parameters for object ($\alpha_{\rm o}$) and probe ($\alpha_{\rm p}$) in ePIE and LoRePIE algorithms.